# TPK: Trustworthy Trajectory Prediction Integrating Prior Knowledge For Interpretability and Kinematic Feasibility

### v2

@@ I Introduction
While neural networks have significantly advanced trajectory prediction by learning complex patterns from large datasets [2], their data-driven nature can lead to illogical or flawed predictions that deviate from human reasoning. For example, as illustrated in Fig. 1(a), state-of-the-art models may predict trajectories influenced by high importance assigned to agents that are irrelevant to the scenario. These limitations undermine the trustworthiness of trajectory prediction systems, especially in rare but safety-critical scenarios, ultimately hindering the practical application of these methods in autonomous driving [3].
@@
Rather than solely pursuing further improvements in the already high accuracy of state-of-the-art trajectory prediction models, we follow the argumentation of [3] that enhancing their trustworthiness is the critical next step. To achieve this, we propose a novel interaction prior applicable to all agent classes, compare two integration methods, and guarantee kinematic feasibility specific to each agent type. Integrating the interaction prior into agent-to-agent attention aligns the model prediction with human intuition, enhancing its interpretability, as shown in Fig. 1(b). Additionally, enforcing kinematic feasibility through agent-specific kinematic layers ensures predictions adhere to physical laws. Overall, we introduce three key contributions to enhance the trustworthiness of trajectory prediction networks:
@@ III Methodology
This section details our approach to improving the trustworthiness and interpretability of trajectory prediction networks, as illustrated in Fig. 2. We leverage HPTR [6] as our backbone, a state-of-the-art transformer-based network known for its efficiency and extensibility. To address the limitations highlighted in  Section II, we propose enhancing encoder interpretability by integrating a novel interaction prior applicable to all agent classes and improving decoder feasibility through a kinematic model, with special emphasis on a novel pedestrian-specific kinematic model.
@@ III-A2 Interaction Prior
Firstly, when agent $j$ enters the focal agent $i$’s personal space, it becomes crucial for the focal agent to monitor $j$’s actions in order to anticipate potential interactions. Eq. 1 defines the interaction importance $\beta_{ij}^{\,A}$ assigned from the focal agent $i$ to agent $j$. Here, $r_{i}=[x_{i},y_{i}]^{T}$ and $v_{i}=[v_{x_{i}},v_{y_{i}}]^{T}$ represent the 2D position and velocity of agent $i$ at the last observation time $T_{o}$, respectively.
@@
Secondly, if the focal agent intrudes into agent $j$’s personal space, it actively triggers an interaction with $j$. In this case, the focal agent monitors whether $j$ reacts according to its expectation. Therefore, we assess how the strength of $j$’s potential at the position of $i$ changes over time as $i$ and $j$ move. This aspect of the interaction is formalized as $\beta_{ij}^{\,B}$ in Eq. 2, where $r_{i}^{*}=r_{i}+v_{i}\cdot N_{DG}\,\Delta t$ represent the estimated future position of agent $i$, with $\Delta t$ denoting the time interval between state observations.
@@ III-A3 Interaction Prior Integration
To the best of our knowledge, no previous work has incorporated an interaction prior into an attention mechanism for mixed traffic trajectory prediction or adequately addressed its implications. Thus, inspired by integration methods outlined in Section II-A, we investigate two promising approaches to compare their impact on the attention scores’ interpretability.
@@
MnR is a simple and easy-to-integrate approach. It combines the network’s predicted attention scores $\alpha_{ij}^{\text{pred}}$ with the prior interaction importance scores $\beta_{ij}$ by multiplying them. The combined attention score $\alpha_{ij}^{\text{cmb}}$ for neighbor $j$ of agent $i$ is computed as per Eq. 3. We view $\alpha_{ij}^{\text{cmb}}$ as a trustworthy attention score guided by the prior. Consequently, the attention layer uses it in place of $\alpha_{ij}^{\text{pred}}$ to fuse the embeddings of surrounding agents $J_{i}$ with the focal agent’s embedding.
@@
GnL is a more flexible approach for prior integration, inspired by the success of gating mechanisms in interpreting feature importance in NLP and computer vision [15, 19]. It introduces a gating multi-layer perceptron (MLP) to compute gating values for agent $i$, based on the agent embedding $u_{i}$, the embeddings of its $K$-nearest neighbors $u_{j}$, predicted attention scores $\alpha_{i}^{\text{pred}}\in[0,1]^{|J_{i}|}$ and the prior’s interaction importance scores $\beta_{i}\in[0,1]^{|J_{i}|}$, as shown in Eq. 4. In this formulation, sigm is a sigmoid function, $W\in\mathbb{R}^{|J_{i}|\times\text{hidden}}$ is the weight matrix, and $b\in\mathbb{R}^{|J_{i}|}$ is the bias vector of the gating MLP.
@@
Subsequently, the combined attention score between agents $i$ and $j$ is computed as a weighted sum of the predicted attention and the prior, as shown in Eq. 5. The learned gating value $\sigma_{ij}$ determines the extent to which the prior influences the attention layer’s score. Finally, the combined attention scores are normalized across all neighbors $j$.
@@
To complement the gating mechanism, we introduce an auxiliary loss that minimizes the Kullback–Leibler divergence between the combined attention scores $\alpha_{i}^{\text{cmb}}$ and the prior scores $\beta_{i}$, as illustrated in Eq. 6. The proposed auxiliary loss improves training stability and prevents the network from defaulting to a gating value $\sigma_{ij}$ of 1, which would prioritize predicted attention scores and completely disregard the prior. For multi-head attention, we compute the average of this loss over all attention heads. This allows for variation in the heads while ensuring that they adhere to the prior when their results are put together.
@@ III-B Feasible Decoder
After encoding all agents and map elements with the proposed interpretable encoder, the network utilizes an anchor-based feasible decoder to predict multimodal trajectories for each focal agent $i$ in the scene, followed by our prediction and confidence heads. To ensure kinematic feasibility, we employ action-space predictions with class-specific kinematic layers, following insights from Section II-B. Separate prediction heads for each agent class enable the design of kinematic layers with distinct control inputs and feasibility constraints. Additionally, class-specific classification heads are introduced to give the network greater flexibility in distinguishing between agent classes and classifying their behaviors more independently.
@@ III-B1 Kinematic Priors
For pedestrians, there is no consensus in the literature on a standard kinematic model. Therefore, we compare two state-of-the-art models, the single integrator and the unicycle, both defined as in Trajectron++ [21]. Additionally, we introduce a novel double integrator model, where the control inputs are $x$- and $y$-axis accelerations. An Euler integrator updates the agent’s velocity and position at each time step based on these control inputs. As shown in Fig. 3, the unicycle model overly restricts pedestrian motion, limiting curvature in an unrealistic manner. In contrast, the single integrator model is highly flexible but fails to account for acceleration limits. We find that the double integrator model provides the best trade-off between realism and constraints by allowing pedestrians to flexibly change heading while constraining velocity based on acceleration limits.
@@ III-B2 Kinematic Prior Integration
We derive the control input limits based on physical constraints and empirical observations of vehicle and pedestrian behavior. For vehicles, we adopt the well-established limits of $[-8,8]$ m/s2 for acceleration and $[-0.3,0.3]$ 1/m for curvature [5, 8]. For pedestrians, as prior studies lack specific limits, we draw on research into human walking and acceleration behavior [30, 31, 32] to define appropriate constraints. These studies suggest that a reasonable maximum acceleration for pedestrians is $[-8,8]$ m/s2 while their speed is limited to $[0,10]$ m/s.
@@ IV-C Evaluation Metrics
We assess our approach across three metric categories: accuracy, interpretability, and feasibility. For accuracy, we rely on the popular benchmark metrics in [28, 27]. For interpretability, we measure the difference between the prior’s interaction importance score and the network’s attention scores $\Delta\alpha(i)=\frac{1}{|J_{i}|}\sum_{j\in J_{i}}|\alpha_{ij}-\beta_{ij}|$. We then evaluate the correlation of minADE and $\Delta\alpha$. Depending on the prior integration method, we use either the predicted scores $\alpha^{\text{pred}}_{ij}$ or the combined scores $\alpha^{\text{cmb}}_{ij}$ as the attention scores, selecting the one with the stronger correlation. For feasibility, we evaluate the percentage of trajectory steps that violate feasibility limits specified in Section III-B2. Finally, we calculate class-specific metrics by restricting the evaluation to predictions for a particular class of agents.
@@ V-A Reasoning
In Table I, we investigate whether an integrated interaction prior improves the interpretability of the network’s reasoning and compare various priors and integration methods. Particularly when comparing HPTRi attention scores to DG-SFM’s prior scores, we observe a statistically significant correlation despite the network not being exposed to the prior during training. This correlation suggests a tendency for the network’s prediction to be incorrect when the network’s attention deviates significantly from the prior model’s importance score. This tendency substantiates that aligning the encoder’s reasoning with human reasoning can enhance the prediction’s interpretability and accuracy. Thus, we conclude that minimizing the discrepancy between the model’s attention and the prior is valuable. As larger differences coincide with incorrect predictions, guiding the network’s attention using an interaction prior could improve the prediction.
@@
Comparing the prior models, both SKGACN and DG-SFM lead to a similarly strong correlation. On the other hand, integrating the simple L2 prior does not yield any meaningful correlation. The L2 prior assigns each agent an interaction importance score inversely proportional to its Euclidean distance from the focal agent. Since the L2 prior does not adequately represent the human intuition behind interaction importance, it does not facilitate the same correlation as the SKGACN and DG-SFM priors. DG-SFM resulted in a significantly stronger correlation when no prior is integrated. Furthermore, the mathematically substantiated advantages of its modeling, highlighted in Section III-A2, make it the more robust choice for complex scenarios. When DG-SFM guides our network’s interaction attention, it fits human intuition on qualitative examples like Fig. 1(b), where HPTR failed. Hence, we choose DG-SFM over SKGACN for our final model.
@@ V-B1 Infeasible Motions
Table II: Infeasibility rate in the ground truth data of AV2’s validation split, the trajectories predicted by HPTR [6], and our final model. We deem a trajectory step infeasible if its acceleration or curvature exceeds the class-specific physical limits in Section III-B2. We deem a trajectory infeasible if any of its steps are infeasible.
@@
Table II highlights the presence of significant physically infeasible steps in both the dataset and HPTR’s predictions. Further discussion of the noise and infeasibility in AV2 can be found in [34]. This finding substantiates the rationale behind introducing our kinematic layer, which guarantees physical feasibility for every prediction. Without a kinematic layer, the network likely learns to reproduce these infeasible trajectories, effectively modeling sensor noise to minimize the distance to ground-truth trajectories. Enforcing kinematic constraints ensures that the network produces realistic, kinematically feasible predictions. This not only makes the predictions more trustworthy but also improves the model’s generalizability.
@@
Model Kinematic Layer Class-Specific Interaction Layers Interaction Prior Integration All Agents Vehicle Pedestrian Cyclist
Brier- minFDE6 mAP ($\uparrow$) minADE6 minFDE6 minADE6 minFDE6 minADE6 minFDE6 minADE6 minFDE6
@@
Model Brier- minFDE6 mAP ($\uparrow$) minFDE6 minFDE1 MR2 minADE6 minADE1 Infeasible Steps Infeasible Predictions $\rho\,$(minADE6, $\Delta\alpha$) ($\uparrow$)
@@ V-B2 Accuracy of Kinematic Models
The infeasible ground-truth trajectories in the dataset imply that a network enforcing physical feasibility cannot precisely reproduce some trajectories when adhering to physical limits. For example, the dataset contains trajectories with accelerations as high as 202 m/s2 for vehicles and 31 m/s2 for pedestrians, far exceeding physical limits. To address this, we assess how accurately our kinematic models can approximate the ground-truth trajectories, quantifying their reproduction error. In Table III, we measure this error over all agent classes when using the unicycle model for vehicles and cyclists and the double integrator model for pedestrians, as defined in Section III-B.
The presence of infeasible trajectories in the dataset establishes the reproduction error, discussed in Table III, as the lower bound on prediction accuracy for networks adhering to kinematic limits. While networks without feasibility constraints may achieve higher metric scores by learning noisy or unrealistic behaviors, these predictions are less trustworthy and physically implausible. In contrast, our kinematic layer prevents the model from overfitting to noise, ensuring physically realistic and reliable predictions. We argue that a slight reduction in accuracy in exchange for improved feasibility and trustworthiness is a worthwhile trade-off, making the model more suitable for real-world deployment. Furthermore, we compare the reproduction errors introduced by the three kinematic models for pedestrians, as presented in Table III. Among these, the double integrator model achieves the best balance between physical feasibility and versatility. It is more physically realistic than the single integrator, which lacks the ability to capture acceleration dynamics accurately. At the same time, unlike the unicycle model, the double integrator is flexible enough to reproduce nearly all of the dataset’s trajectories without overly constraining the heading rate. This combination of feasibility and adaptability makes the double integrator model a robust choice, leading to its selection as a pedestrian model in our final network.

### v3

@@ I Introduction
While neural networks have significantly advanced trajectory prediction by learning complex patterns from large datasets [2], their data-driven nature can lead to illogical or flawed predictions that deviate from human reasoning. For example, as illustrated in Fig.˜1(a), state-of-the-art models may predict trajectories influenced by high importance assigned to agents that are irrelevant to the scenario. These limitations undermine the trustworthiness of trajectory prediction systems, especially in rare but safety-critical scenarios, ultimately hindering the practical application of these methods in autonomous driving [3].
@@
Rather than solely pursuing further improvements in the already high accuracy of state-of-the-art trajectory prediction models, we follow the argumentation of [3] that enhancing their trustworthiness is the critical next step. To achieve this, we propose a novel interaction prior applicable to all agent classes, compare two integration methods, and guarantee kinematic feasibility specific to each agent type. Integrating the interaction prior into agent-to-agent attention aligns the model prediction with human intuition, enhancing its interpretability, as shown in Fig.˜1(b). Additionally, enforcing kinematic feasibility through agent-specific kinematic layers ensures predictions adhere to physical laws. Overall, we introduce three key contributions to enhance the trustworthiness of trajectory prediction networks:
@@ III Methodology
This section details our approach to improving the trustworthiness and interpretability of trajectory prediction networks, as illustrated in Fig.˜2. We leverage HPTR [6] as our backbone, a state-of-the-art transformer-based network known for its efficiency and extensibility. To address the limitations highlighted in  Section˜II, we propose enhancing encoder interpretability by integrating a novel interaction prior applicable to all agent classes and improving decoder feasibility through a kinematic model, with special emphasis on a novel pedestrian-specific kinematic model.
@@ III-A2 Interaction Prior
Firstly, when agent $j$ enters the focal agent $i$’s personal space, it becomes crucial for the focal agent to monitor $j$’s actions in order to anticipate potential interactions. Eq.˜1 defines the interaction importance $\beta_{ij}^{\,A}$ assigned from the focal agent $i$ to agent $j$. Here, $r_{i}=[x_{i},y_{i}]^{T}$ and $v_{i}=[v_{x_{i}},v_{y_{i}}]^{T}$ represent the 2D position and velocity of agent $i$ at the last observation time $T_{o}$, respectively.
@@
Secondly, if the focal agent intrudes into agent $j$’s personal space, it actively triggers an interaction with $j$. In this case, the focal agent monitors whether $j$ reacts according to its expectation. Therefore, we assess how the strength of $j$’s potential at the position of $i$ changes over time as $i$ and $j$ move. This aspect of the interaction is formalized as $\beta_{ij}^{\,B}$ in Eq.˜2, where $r_{i}^{*}=r_{i}+v_{i}\cdot N_{DG}\,\Delta t$ represent the estimated future position of agent $i$, with $\Delta t$ denoting the time interval between state observations.
@@ III-A3 Interaction Prior Integration
To the best of our knowledge, no previous work has incorporated an interaction prior into an attention mechanism for mixed traffic trajectory prediction or adequately addressed its implications. Thus, inspired by integration methods outlined in Section˜II-A, we investigate two promising approaches to compare their impact on the attention scores’ interpretability.
@@
MnR is a simple and easy-to-integrate approach. It combines the network’s predicted attention scores $\alpha_{ij}^{\text{pred}}$ with the prior interaction importance scores $\beta_{ij}$ by multiplying them. The combined attention score $\alpha_{ij}^{\text{cmb}}$ for neighbor $j$ of agent $i$ is computed as per Eq.˜3. We view $\alpha_{ij}^{\text{cmb}}$ as a trustworthy attention score guided by the prior. Consequently, the attention layer uses it in place of $\alpha_{ij}^{\text{pred}}$ to fuse the embeddings of surrounding agents $J_{i}$ with the focal agent’s embedding.
@@
GnL is a more flexible approach for prior integration, inspired by the success of gating mechanisms in interpreting feature importance in NLP and computer vision [15, 19]. It introduces a gating multi-layer perceptron (MLP) to compute gating values for agent $i$, based on the agent embedding $u_{i}$, the embeddings of its $K$-nearest neighbors $u_{j}$, predicted attention scores $\alpha_{i}^{\text{pred}}\in[0,1]^{|J_{i}|}$ and the prior’s interaction importance scores $\beta_{i}\in[0,1]^{|J_{i}|}$, as shown in Eq.˜4. In this formulation, sigm is a sigmoid function, $W\in\mathbb{R}^{|J_{i}|\times\text{hidden}}$ is the weight matrix, and $b\in\mathbb{R}^{|J_{i}|}$ is the bias vector of the gating MLP.
@@
Subsequently, the combined attention score between agents $i$ and $j$ is computed as a weighted sum of the predicted attention and the prior, as shown in Eq.˜5. The learned gating value $\sigma_{ij}$ determines the extent to which the prior influences the attention layer’s score. Finally, the combined attention scores are normalized across all neighbors $j$.
@@
To complement the gating mechanism, we introduce an auxiliary loss that minimizes the Kullback–Leibler divergence between the combined attention scores $\alpha_{i}^{\text{cmb}}$ and the prior scores $\beta_{i}$, as illustrated in Eq.˜6. The proposed auxiliary loss improves training stability and prevents the network from defaulting to a gating value $\sigma_{ij}$ of 1, which would prioritize predicted attention scores and completely disregard the prior. For multi-head attention, we compute the average of this loss over all attention heads. This allows for variation in the heads while ensuring that they adhere to the prior when their results are put together.
@@ III-B Feasible Decoder
After encoding all agents and map elements with the proposed interpretable encoder, the network utilizes an anchor-based feasible decoder to predict multimodal trajectories for each focal agent $i$ in the scene, followed by our prediction and confidence heads. To ensure kinematic feasibility, we employ action-space predictions with class-specific kinematic layers, following insights from Section˜II-B. Separate prediction heads for each agent class enable the design of kinematic layers with distinct control inputs and feasibility constraints. Additionally, class-specific classification heads are introduced to give the network greater flexibility in distinguishing between agent classes and classifying their behaviors more independently.
@@ III-B1 Kinematic Priors
For pedestrians, there is no consensus in the literature on a standard kinematic model. Therefore, we compare two state-of-the-art models, the single integrator and the unicycle, both defined as in Trajectron++ [21]. Additionally, we introduce a novel double integrator model, where the control inputs are $x$- and $y$-axis accelerations. An Euler integrator updates the agent’s velocity and position at each time step based on these control inputs. As shown in Fig.˜3, the unicycle model overly restricts pedestrian motion, limiting curvature in an unrealistic manner. In contrast, the single integrator model is highly flexible but fails to account for acceleration limits. We find that the double integrator model provides the best trade-off between realism and constraints by allowing pedestrians to flexibly change heading while constraining velocity based on acceleration limits.
@@ III-B2 Kinematic Prior Integration
We derive the control input limits based on physical constraints and empirical observations of vehicle and pedestrian behavior. For vehicles, we adopt the well-established limits of [8,8][-8,8][ - 8 , 8 ] m/s2 for acceleration and [0.3,0.3][-0.3,0.3][ - 0.3 , 0.3 ] 1/m for curvature [5, 8]. For pedestrians, as prior studies lack specific limits, we draw on research into human walking and acceleration behavior [30, 31, 32] to define appropriate constraints. These studies suggest that a reasonable maximum acceleration for pedestrians is [8,8][-8,8][ - 8 , 8 ] m/s2 while their speed is limited to [0,10][0,10][ 0 , 10 ] m/s.
@@ IV-C Evaluation Metrics
We assess our approach across three metric categories: accuracy, interpretability, and feasibility. For accuracy, we rely on the popular benchmark metrics in [28, 27]. For interpretability, we measure the difference between the prior’s interaction importance score and the network’s attention scores $\Delta\alpha(i)=\frac{1}{|J_{i}|}\sum_{j\in J_{i}}|\alpha_{ij}-\beta_{ij}|$. We then evaluate the correlation of minADE and $\Delta\alpha$. Depending on the prior integration method, we use either the predicted scores $\alpha^{\text{pred}}_{ij}$ or the combined scores $\alpha^{\text{cmb}}_{ij}$ as the attention scores, selecting the one with the stronger correlation. For feasibility, we evaluate the percentage of trajectory steps that violate feasibility limits specified in Section˜III-B2. Finally, we calculate class-specific metrics by restricting the evaluation to predictions for a particular class of agents.
@@ V-A Reasoning
In Table˜I, we investigate whether an integrated interaction prior improves the interpretability of the network’s reasoning and compare various priors and integration methods. Particularly when comparing HPTRi attention scores to DG-SFM’s prior scores, we observe a statistically significant correlation despite the network not being exposed to the prior during training. This correlation suggests a tendency for the network’s prediction to be incorrect when the network’s attention deviates significantly from the prior model’s importance score. This tendency substantiates that aligning the encoder’s reasoning with human reasoning can enhance the prediction’s interpretability and accuracy. Thus, we conclude that minimizing the discrepancy between the model’s attention and the prior is valuable. As larger differences coincide with incorrect predictions, guiding the network’s attention using an interaction prior could improve the prediction.
@@
Comparing the prior models, both SKGACN and DG-SFM lead to a similarly strong correlation. On the other hand, integrating the simple L2 prior does not yield any meaningful correlation. The L2 prior assigns each agent an interaction importance score inversely proportional to its Euclidean distance from the focal agent. Since the L2 prior does not adequately represent the human intuition behind interaction importance, it does not facilitate the same correlation as the SKGACN and DG-SFM priors. DG-SFM resulted in a significantly stronger correlation when no prior is integrated. Furthermore, the mathematically substantiated advantages of its modeling, highlighted in Section˜III-A2, make it the more robust choice for complex scenarios. When DG-SFM guides our network’s interaction attention, it fits human intuition on qualitative examples like Fig.˜1(b), where HPTR failed. Hence, we choose DG-SFM over SKGACN for our final model.
@@ V-B1 Infeasible Motions
Table II: Infeasibility rate in the ground truth data of AV2’s validation split, the trajectories predicted by HPTR [6], and our final model. We deem a trajectory step infeasible if its acceleration or curvature exceeds the class-specific physical limits in Section˜III-B2. We deem a trajectory infeasible if any of its steps are infeasible.
@@
Table˜II highlights the presence of significant physically infeasible steps in both the dataset and HPTR’s predictions. Further discussion of the noise and infeasibility in AV2 can be found in [34]. This finding substantiates the rationale behind introducing our kinematic layer, which guarantees physical feasibility for every prediction. Without a kinematic layer, the network likely learns to reproduce these infeasible trajectories, effectively modeling sensor noise to minimize the distance to ground-truth trajectories. Enforcing kinematic constraints ensures that the network produces realistic, kinematically feasible predictions. This not only makes the predictions more trustworthy but also improves the model’s generalizability.
@@
Model
Kinematic
Layer
Class-Specific
Interaction
Layers
Interaction
Prior
Integration
All Agents Vehicle Pedestrian Cyclist
Brier-
minFDE6
mAP ($\uparrow$) minADE6 minFDE6 minADE6 minFDE6 minADE6 minFDE6 minADE6 minFDE6
@@
Model
Brier-
minFDE6
mAP ($\uparrow$) minFDE6 minFDE1 MR2 minADE6 minADE1
Infeasible
Steps
Infeasible
Predictions
$\rho\,$(minADE6, $\Delta\alpha$) ($\uparrow$)
@@ V-B2 Accuracy of Kinematic Models
The infeasible ground-truth trajectories in the dataset imply that a network enforcing physical feasibility cannot precisely reproduce some trajectories when adhering to physical limits. For example, the dataset contains trajectories with accelerations as high as 202 m/s2 for vehicles and 31 m/s2 for pedestrians, far exceeding physical limits. To address this, we assess how accurately our kinematic models can approximate the ground-truth trajectories, quantifying their reproduction error. In Table˜III, we measure this error over all agent classes when using the unicycle model for vehicles and cyclists and the double integrator model for pedestrians, as defined in Section˜III-B.
The presence of infeasible trajectories in the dataset establishes the reproduction error, discussed in Table˜III, as the lower bound on prediction accuracy for networks adhering to kinematic limits. While networks without feasibility constraints may achieve higher metric scores by learning noisy or unrealistic behaviors, these predictions are less trustworthy and physically implausible. In contrast, our kinematic layer prevents the model from overfitting to noise, ensuring physically realistic and reliable predictions. We argue that a slight reduction in accuracy in exchange for improved feasibility and trustworthiness is a worthwhile trade-off, making the model more suitable for real-world deployment. Furthermore, we compare the reproduction errors introduced by the three kinematic models for pedestrians, as presented in Table˜III. Among these, the double integrator model achieves the best balance between physical feasibility and versatility. It is more physically realistic than the single integrator, which lacks the ability to capture acceleration dynamics accurately. At the same time, unlike the unicycle model, the double integrator is flexible enough to reproduce nearly all of the dataset’s trajectories without overly constraining the heading rate. This combination of feasibility and adaptability makes the double integrator model a robust choice, leading to its selection as a pedestrian model in our final network.